North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6.

As the Covid-19 pandemic had reduced the requirement for locomotives on the North Yorkshire Moors Railway, in June 2020 a small group of volunteers, working within the Covid-19 protocol, made an early start on the plan to lift the boiler out of the frames. This was to allow an external examination of the boiler by the insurance company, and to allow the boiler ticket to be extended to 2028 - 10 years from the date of the internal examination. The opportunity was taken at this time to service the axle boxes, horns and related "bottom end" components.

On 27th October the Q6 made 2 light engine test runs to Goathland wth no problems. On Thursday 28th it made another light engine trip to Goathland and then took the 14.30 train to pickering and return with no problems. Many thanks to all those, both NELPG and NYMR, who have worked so hard over the last two years.



Q6 - 22 December 2021

Ian reports a cold fine day at Grosmont on Wednesday with Steve Hyman, Ed Bolam and himself.

Not a lot to report on the Q6, barring being lit with a warming fire during the afternoon with the possibility of working the Grosmont Santa service yesterday. It's previous working was on Monday 20th with Terry Newman driving and Peter James firing (an all NELPG crew!) on the first half of Pickering Santas . The second half was done with Black Five 5428. This was due to transferring empty stock through to Pickering.

As this is his last report for 2021, Ian wishes everyone a happy Christmas and all the best for a healthy and successful 2022.

 

Q6- 15 December 2021

Ian reports a fine sunny day at Grosmont last Wednesday, with only Steve Hyman, Ed Bolam and Ian Pearson present.

After the morning cuppa, two pallets were removed from the MPD truck with the JCB, one of brake blocks and one containing a few rail chairs, fishplates and a bag of rail keys. These had been collected by the MPD's Sue Smeaton from Hopetown as part of the clear out of the building. All the brake blocks have been stored under the buffer beam at the back of No 8 Road in Deviation Shed. The 6 Chairs and fishplates are at the side of Armstrong Oilers. The tally of brake blocks is 7 new BR2, 9 partly worn BR2, 1 partly worn BR4 and 2 new BR4 - 19 in all. The two NELPG rail trolleys were moved out of Deviation Shed and put near Armstrong Oilers ready to load onto the MPD truck on Monday morning which is going back to Hopetown to pick up some more rail chairs. We are most grateful for the NYMR's assistance with this work.

Steve Hyman brought in a new four step ladder which he had acquired.

This was inspected by MPD fitter Duncan Jackson and tagged and recorded (see photo).

New four step ladder from Steve Hyman for Deviation Shed, inspected, tagged and recorded on 15 December 202 - Ian Pearson

Chris Henwood has been employed on Q6 repairs this week, with repacking the left and right valve gland packings. In removing the righthand piston packing, some bright marks were identified at top and bottom of the packing so these were refitted. Tests were carried out to see if the new lubricator was delivering oil to the axle boxes. Oil was evident when the pipe on top of the box was removed and the pump manually turned. The righthand injector steam valve was taken apart and ground in, then refitted. Three more rubber pipes and jubilee clips were connected to the horn guide delivery pipes. A warming fire was lit during the afternoon in preparation for a steam test the next day.

63395 as Rudolph the red nosed reindeer, ready for NYMR Santa specials on 15 December 2021 - Steve Hyman

Although the Q6 didn’t work last Saturday, it was in steam on standby but worked as Rudolph (see photo) on Sunday on Grosmont Santas, top and tailed with the Class 37.

 

Q6 - 7/8 December 2021

Ian reports that on Tuesday he called in at Grosmont to find the MPD crew in the process of removing the J27 cab.  While this was going on, he and Nik Proctor cleared the pit of ash. Ian then put the pump on inside Deviation shed as the pit in there was quite deep. During the afternoon Ian put a warming fire in the Q6 ready for a steam test on the next day and ended up staying until 4pm.

On Wednesday, Ian was on his own until Jon arrived just before lunchtime. A mega shunt was going on at Deviation Shed, as a result of which Vera Lynn has now been moved outside and the well wagon with the generator on is now on No 6 Road in Deviation shed. The J27 is at the back of the Shed on No 7 Road with the Class 24 Helen Turner in front. The Q6 was in steam and work was being carried out on the lefthand side piston packing by Chris Henwood. The righthand cylinder cosmetic cover had been removed to investigate a leak from the main cover that we already knew about, so the cosmetic cover was refitted by Ian and Jon. Later in the afternoon, after Chris had refitted the piston packing, the Q6 was given a run up the head shunt in readiness for its outing on today's diner.

Ian is planning to go to Grosmont tomorrow (Friday) with Robson Hewitson to remove some items from the J27 to allow us to cut out a section of cladding in the area where the broken stays are situated.

A supplementary report from Ian Pearson. On Friday the Q6, a union, which is labelled L1, was leaking oil. It was found to be very loose and was tightened up and tested by turning the ratchet wheel - the leak had been cured. Robson and Ian then had to go and give the 9F a test run up the head shunt.

 

Q6 - 1 December 2021

Ian Pearson reports that Wednesday was a cold day with patchy clouds, some sunshine but heavy hail showers. Present were David Potter, Ed Bolam, Paul Jameson and Ian. Bill Dobson is likely to be away in Scotland for most of December.

After the Q6 was shunted on to the pit near Deviation shed, some work left over from Tuesday was carried out. The regulator gland had been leaking: this was tightened up, along with the steam reverser cataract cylinder rear gland. A bolt which secures the steam valve spindle to the operating rod that is attached to the steam reverser was found to be fouling the rod's movement and not allowing steam to be shut off from the valve on top of steam cylinder. When opened it allows oil to drain into the cylinder small slide valve. This bolt was replaced with a shorter one.

The fire grate was cleaned and the ash pan cleared out and a new warming fire set and lit later in the afternoon as the Q6 was booked to work the Moorlander diner on Thursday lunchtime. Attached is a photo Ian took on the Thursday of the Q6, masquerading as Prancer, having its smokebox emptied by fireman Mike Burn after its trip to Pickering and back with the Moorlander. It is expected to be the regular locomotive for the diner throughout December.

63395 having its smokebox emptied by fireman Mike Burn after its round trip to Pickering with the Moorlander diner on 1 December 2021 - Ian Pearson

The Defibrillator was checked and was OK.

 

Q6 - 24/25 November 2021

Ian Pearson reports a cold fine day at Grosmont on Wednesday, but with just David Potter and Jon Bradley joining Ian. Bill had had to rush off to Scotland at short notice for family business.

David worked on the axle box lubricator all day and completed the rebuild. L1 pump had the original thimble installed as the new one would not fit. The lubricator was tested satisfactorily and delivered oil from all eight oil lines. Jon continued to work around the new gate area, fitting concrete slabs and building up the step at the front of the access route. There is still more work for Jon to do in this area next time. Ian fitted the pressure gauges in the cab. The metal plate that Steve Hyman made last week had a stabilisation bar welded to the underneath side and then was fitted around the lefthand tender water valve. The tender to loco water hoses were connected by the shed staff.

Thursday saw Ian at Grosmont again, but on a cold slightly drizzly morning. The Q6 had been shunted outside the MPD on No 5 Road with a hose connected to the injector overflow pipe filling the boiler. He then set the fire. After the boiler had sufficient water, the hose was transferred to the tender. The warming fire was then lit. As Lucie and No 29 were inside the shed on No 5 Road and were ready for moving outside, the Q6 was shunted round to the water column and filled. It was then moved out of the way at the east side of Deviation shed on the wheel drop road. Ian managed a couple of jobs lagging the copper pipe that feeds the ash pan spray. He then fitted a couple of jubilee clips to the rubber extension pipes that feed oil to the horn guides. Later, with the warming fire burning nicely, the Q6 was returned outside the MPD on No 5 Road ready for a steam test today (Friday 26th).

 

Q6 - 17 November 2021

Ian reports a fine mild day at Grosmont on Wdnesday with Bill Dobson, Steve Hyman, Ed Bolam, Nigel Hall, Nigel Bill, Jon Bradley, Gordon Wells, David Potter and Ian Pearson.

That morning, the Q6 was moved from No 2 road onto the pit outside Deviation Shed. This allowed Ian Pearson to work underneath to adjust the oil feed pipes into a position where the oil drops centrally along the top of the horn guides. Bill Dobson and David Potter pressure tested the atomisers and found them to be OK with no leakage from joints and castings. These are intended to be fitted in the New Year. Afterwards they concentrated on the new axle box lubricator, changing the thimbles, eight of which were made at Hopetown last week. Bit of a fiddly job, and only four were fitted this week.

Nigel Bill arrived with a new battery for the defibrillator which was fitted and is now functioning again. Nigel then set about sorting out and tidying up our cabinet drawers upstairs, resulting in some rubbish being thrown out. Steve Hyman is in the process of making a plate around the lefthand tender water valve. This has been cut to size and a stability bracket was made but it requires welding under the new plate next week. Gordon Wells came in and repaired the outstanding PAT testing failures. These were sent to Hopetown for re-testing on Thursday 18th.

Jon Bradley arrived and finished the new access gate, but it does need a step fitting to the front.

Later, Nigel Hall and Ed Bolam lagged the steam heat pipe after an earlier complaint from a fireman who had nearly received a burn while disposing of the ashes from the ash pan. The pipe area lagged was from the front damper forwards about 6 feet.

In the afternoon, David Potter had his site safety induction with Paul Whickham.

 

Q6 - 10 November 2021

Ian reports a mild overcast Wednesday at Grosmont this week, with Bill Dobson, Steve Hyman, Ed Bolam, Nigel Hall, Terry Newman, David Potter and Ian Pearson present.

On the Q6, Bill Dobson and David Potter serviced and cleaned the Q6 whistle, then refitted it back onto the boiler. They then started working on the new axle box lubricators, removing one of the thimbles. This thimble measures the amount of the pump stroke when delivering oil to the axle boxes. Bill would like the lubricator to deliver more oil, so he is sending one of the thimbles as a pattern to Hopetown with Ed Bolam for Arthur Jenkins to make eight new ones, slightly shorter, so they will deliver more oil. Later they began to hydraulic test the atomisers, but that was not successful at the moment as the fittings have different threads so some new pipe work is required.

Ian Pearson and Steve Hyman patched up the Q6 back sheet with Goliath tape then fitted it to the cab, so it will hopefully keep the inside of the cab dry if it rains. It may be necessary though to acquire a new one in the New Year, but we will have to see how the patchwork repairs hold during the Christmas running. Steve also started to make a cover that fits around the tender water shut off linkage. He will continue this next week. Ed Bolam has measured up for two new securing pins to be made for the Q6 front windows. Hopefully, Arthur Jenkins will make these at Hopetown.

Ian has spoken to Nick Simpson and Paul Wickham regarding an induction course for new volunteer David Potter. After discussing with Paul, he will take David on an induction course next Wednesday 17 November. David is aware.

 

Q6 - 2/3 November 2021

Ian Pearson reports a cool fine day with some bubbly clouds at Grosmont on Tuesday with only Ian and Bill Dobson to start with, although Bill could not stay for long due to other commitments with a horse. As Bill was leaving, Gordon Wells arrived. Then around lunchtime Jon Bradley clocked on, so Ian wasn’t Billy no mates!

After the PAT testing from last Thursday with Nigel Bill in attendance, there were some electrical repairs to carry out to which Gordon attended, but leaving the oil filled radiator to repair next time, as it needs a new cable and plug. Ian put all the tested equipment that had passed, back in the workshop in their respective places. The Defibrillator, which was checked out last Thursday and all was OK, started giving a warning bleep in the morning, and showing a red light indicating the battery was ready for replacement. Nigel Bill was informed and Richard Wheeler is ordering a replacement. The replacement stove chimney extension had been fitted sometime over the weekend by Andy Whiffin, but, as it was raining when he was on site, he could not repair the cracks in the two roofing sheets which were leaking. He will be back later to deal with them.

Jon spent all afternoon removing some of the old fence and fitting a new gate to give a safer access to Deviation Shed. This critical path can now be used to give access from the MPD, although there is still some finishing off to do which Jon will complete shortly. The old access, direct on to the ash pit running line, will be kept to allow heavy items to be moved into Deviation Shed when required, but will be kept locked when not needed.

Wednesday, in contrast, was a day of rain and hail showers, some of which were very heavy. Six were present to brave the rain - Bill Dobson, Steve Hyman, Ed Bolam, Nigel Hall, David Potter and Ian Pearson himself.

It turned into a joint J27/Q6 working party and on the Q6, Bill and David removed the whistle valve for examination and gave it to Owain for machining. They then turned their attention to the lefthand big end, which was found to be a lot warmer than the righthand side during last week's running. The large cotter was slightly eased and a new securing cotter made and fitted. Ian fitted 16 new trimmings to the horn guide brass oil boxes.

With the operating season now over until Christmas, and both locomotives only recently returned to traffic, it has been decided to reduce our working parties to one day a week for the moment. So, the next working party will be on Wednesday 10 November.

 

23rd to 28th October 2021 - Q6 Test Runs

The end of the line! Bill Dobson takes a bow after the Q6 hauls it's first train since overhaul - Jonathan Bradley

The Q6 is finally back in revenue earning service, so many thanks to Bill and Ian Pearson in particular, but also to everyone else who has been involved over the past two years, for all their hard work in overcoming frustrating setbacks and problems, culminating in the successful arrival at Pickering on the Thursday of 63395 with the 14.30 departure from Grosmont.

Read more...
 

Q6 - 19/20 October 2021

A wet but very mild day for October on Tuesday at Grosmont for Bill Dobson, Gordon Wells and Ian Pearson. As Ian was on his own for a while at the start of the day, he set up a table in Deviation Shed to put out all our electrical equipment that requires PAT testing, ready for when the electrician comes. As he was about to start, Caron Webster (who is apparently leaving as Head of NYMR Finance), came in looking for a job, so she moved most of our equipment onto the table and later cleaned all the grease and grime from the righthand cylinder copper ring. Thanks to Caron for that contribution.

Bill had been notified by Team Industrial Services that they would be machining the face of the cylinder on Saturday 23rd, so Ian then started to remove the twenty cylinder cover studs. They took some effort to remove but, with some help from Gordon and Chris Henwood, eventually eighteen were successfully removed. But there were two which even the strong arm shed fitters couldn’t remove - a job for the next day. Andy Whiffin, the local building contractor, also called to check out the fitting of our stove chimney extension, and also the sealing of some leaks in Deviation Shed roof. He plans to come in on Saturday morning (23rd) to carry out the work . Bill Dobson, Ian Pearson and Ed Bolam will be present for both.

On Wednesday, there was a surfeit of riches, with Bill Dobson, Steve Hyman, Ed Bolam, Nigel Hall, Nigel Bill, Chris Lawson and Ian Pearson all in attendance on a cold but fine morning, which turned into heavy rain around lunchtime.We also had a visitor, David Potter, a railway director and professional engineer, who was interested in getting involved and helping the Group in some way. As a result of conversations with us, he intends to become a member and hopes to come back to Grosmont for next Wednesday's working party with his overalls.

All our ladders and steps now come under the MPD inspection system, the latest of which was carried out during the day. We did have five sets of ladders when it started, but, after the inspection, three sets have been scrapped due to various faults. We also had two sets of steps, one of which has also had to be scrapped. All were cut in half and put into the skip. Please note when using ladders in future, make sure they are safe and have the current colour tag fitted - it is shown on the back of the workshop door.

On the Q6, both cylinder studs left in from Tuesday were removed by MPD staff - nuts had to be welded on to remove them. All the cylinder studs were then cleaned by Ed Bolam and Steve Hyman and a die nut run down each one. The tender tank was filled with water and last week's fire remains removed and ash pan cleaned out. In the cab, the wooden cover over the lefthand sandbox lid was varnished by Steve Hyman. The copper ring from the cylinder was annealed.

Just as the rain started, the locomotive was shunted onto the pit so some modifications could be made to the horn guide pipework by adding a piece of flexible pipe to allow oil to be accurately delivered in spite of the up and down movement of the locomotive, and without the more rigid copper pipe being damaged. Only two pipes were successfully modified in the pouring rain, which soaked all concerned, before No 29 needed to be on the pit, and the Q6 had to be put back into the running shed. Hopefully this work will be completed next week when the weather might also be better.

 
  • «
  •  Start 
  •  Prev 
  •  1 
  •  2 
  •  3 
  •  4 
  •  5 
  •  6 
  •  7 
  •  8 
  •  9 
  •  10 
  •  Next 
  •  End 
  • »


Page 1 of 35